OpenStudy (anonymous):

How do you simplify this equation into a value for y ? y = (Sqrt[2]/2) / E^((Sqrt[2]/2)^2) ​

OpenStudy (usukidoll):

oh wow what is this lol

OpenStudy (anonymous):

y = (Sqrt[2]/2) / E^((Sqrt[2]/2)^2) I think it's this.. ​\[y =\frac{ \frac{\sqrt{2}}{2} }{ E^{\frac{\sqrt{2}}{2}^2} ​}\]

OpenStudy (anonymous):

Im supposed to be able to do this by hand... :/

OpenStudy (usukidoll):

OMG who assigns this?

OpenStudy (anonymous):

its the worst calculus course ever invented.. bill davis, horacio porta and Jerry Uhl

OpenStudy (anonymous):

apparently a flipped curriculum.. where they explain nothing, and try to have the students work it out for themselves.

OpenStudy (usukidoll):

ffffffffffffffffffffffff

OpenStudy (usukidoll):

\[(\frac{\sqrt{2}}{2})^2\] I would start expanding this first.

OpenStudy (usukidoll):

how is this Calculus? More like torture.

OpenStudy (anonymous):

maybe \[ (\frac{2^.5} { 2 }) (\frac{2^.5}{ 2 })\] \[ \frac{(2^.5)(2^.5)} { 4 } \]

OpenStudy (anonymous):

Ive been thinking the same thing.. lol

OpenStudy (usukidoll):

ummm ... \[\frac{\sqrt{2}}{2}(\frac{\sqrt{2}}{2}) = \frac{2}{4} \rightarrow \frac{1}{2}\]

OpenStudy (anonymous):

woah.. is that true?

OpenStudy (anonymous):

yes it is

OpenStudy (usukidoll):

\[\LARGE e^{\frac{1}{2}}\]

OpenStudy (usukidoll):

darn it ... I sort of forgot some of the e rules I know for \[\large e^{2+x} \rightarrow e^2e^x\]

OpenStudy (anonymous):

so I think I can move the 2 in the numerator down to the denominator

OpenStudy (usukidoll):

I need to relook it up... I think it involves one of em .

OpenStudy (anonymous):

yes thats true..

OpenStudy (usukidoll):

can't remember =/

myininaya (myininaya):

hey so you are suppose to "simplify" the expression called y? and y is given as: \[y=\frac{\frac{\sqrt{2}}{2}}{e^{(\frac{\sqrt{2}}{2})^2}}\]

OpenStudy (usukidoll):

yeah and I already expanded the bottom

OpenStudy (usukidoll):

because that's an easy place to start

OpenStudy (anonymous):

so I think we get y= 2^(1/2) / 2 E^(1/2)

OpenStudy (usukidoll):

where did that 2 come from? ^ at the denominator

OpenStudy (anonymous):

btw is there an easy way to enter these equations? that equation tool takes me forever.

myininaya (myininaya):

you can also write that as: \[y=\frac{\sqrt{2}}{2} \cdot \frac{1}{e^\frac{1}{2}}\]

myininaya (myininaya):

you can also chose to multiply the top and bottom by e^(1/2) to rationalize the denominator

myininaya (myininaya):

well sorta rationalize anyways since e isn't rational :p

OpenStudy (usukidoll):

\[\large y=\frac{\sqrt{2}}{2} \cdot \frac{1}{e^\frac{1}{2}} \] I rather have this XD can't we flip the second fraction over or use one of the exponent laws for that 1/e^{something}

OpenStudy (anonymous):

I cant believe these questions aren't all over open study.. I must be the only idiot in the world doing this course..

OpenStudy (usukidoll):

cough common core cough xD

myininaya (myininaya):

yes you can do this: \[y=\frac{\sqrt{2}}{2 e^\frac{1}{2}} \cdot \frac{e^\frac{1}{2}}{e^\frac{1}{2}}\]

OpenStudy (usukidoll):

wasn't there also an ... I forgot what's it's called but if I have something like this... I can split it up...what's it called? \[\large e^{2+x} \rightarrow e^2e^x \]

myininaya (myininaya):

that is one the law of exponents

OpenStudy (anonymous):

I have this in notes... The main laws of exponents are e^(a + b) = e^a e^b e^(a - b) = e^a/e^b (e^a)^b = e^(a b) . a / E^n = a E^-n negative exponent law

OpenStudy (usukidoll):

ah... so would this work? \[\large e^{\frac{1}{2}} \rightarrow e^{1}e^{-2}\]

myininaya (myininaya):

no

myininaya (myininaya):

\[e^1e^{-2}=e^{1-2}=e^{-1}=\frac{1}{e}\]

OpenStudy (usukidoll):

oh xD now I remember

myininaya (myininaya):

have you considered getting rid of the square root part on bottom as I suggested earlier ?

OpenStudy (usukidoll):

e and the other exponent laws are different \[n^{a-b} = \frac{n^a}{n^b}\]

myininaya (myininaya):

n can be e

OpenStudy (usukidoll):

yeah \[\frac{\sqrt{2}}{2}(\frac{\sqrt{2}}{2}) = \frac{2}{4} \rightarrow \frac{1}{2} \]

OpenStudy (usukidoll):

oh whoops yeah for something else not the exponent law thing

OpenStudy (anonymous):

maybe log.. can be applied?

myininaya (myininaya):

remember this: \[y=\frac{\sqrt{2}}{2 e^\frac{1}{2}} \cdot \frac{e^\frac{1}{2}}{e^\frac{1}{2}} \\ y= \frac{\sqrt{2} e^\frac{1}{2}}{2 e}=\frac{\sqrt{2 e}}{2e}\]

myininaya (myininaya):

anyways that is the form I would leave it in but you may think the square thing looks better on bottom and if you think that you can multiply top and bottom by sqrt(2e) and simplify from there but this form above is the one I would leave it as

myininaya (myininaya):

square root thing*

OpenStudy (anonymous):

My calculator says this will end up being 1/Sqrt[2E] and I would get a value of 0.428882

myininaya (myininaya):

ok if you want that other result as I said you can put the square root thing in the bottom by multiplying top and bottom by sqrt(2e) and simplifying from there

OpenStudy (anonymous):

are we on track?

myininaya (myininaya):

\[y=\frac{\sqrt{2e}}{2e} \\ \text{ this is my preferred form } \text{ multiply } \sqrt{2e} \text{ on bottom and \top and simplify from } \\ \text{ there if you want } y=\frac{1}{\sqrt{2e}}\]

OpenStudy (dan815):

sounds like a fun course :)

OpenStudy (anonymous):

lol, dan you would own this course.. :)

OpenStudy (anonymous):

is there a trick to turning this into a value that I can set into a chart.. I'm supposed to plot points with this, by hand.

myininaya (myininaya):

example: \[\frac{\sqrt{u}}{u} \\ \frac{\sqrt{u}}{u} \cdot \frac{\sqrt{u}}{\sqrt{u}} =\frac{u}{u \sqrt{u}}=\frac{1}{\sqrt{u}}\]

OpenStudy (anonymous):

ah I like that perspective.. thnx. myininaya, I will add that to the list..

OpenStudy (usukidoll):

oh I see.. after we got \[e^\frac{1}{2}\] we changed it in radical form and rationalize!

OpenStudy (dan815):